How do you write an expression for a polynomial of degree 3 with zeros, x= -3,1,-1 and f(2)=15?
Answer by stanbon(75887)

How do you write an expression for a polynomial of degree 3 with zeros, x= -3,1,-1 and f(2)=15?
----------
Since you have those three zeroes you know the expression
has factors (x+3)(x-1)(x+1); but it may also have a constant
multiplier; so you write:
f(x)=a(x+3)(x-1)(x+1)
f(2)=a(5)(1)(3)=15
15a=15
a=1
So the equation is f(x)=1(x-3)(x-1)(x+1)
Cheers,
Stan H.